# Simulation Based Bayesian Optimization

### v2

@@ 1 Introduction
Bayesian Optimization (BO, [21]) stands as an efficient approach for identifying the extrema of a black-box objective function. Its effectiveness originates from its capacity to integrate prior knowledge about the function with the evaluations previously conducted, enabling informed decisions on where to carry out additional evaluations. Bayesian Optimization involves the construction of a Bayesian surrogate model for $f(\boldsymbol{x})$. Subsequently, the next evaluation point is selected through the maximization of an acquisition function [9].
@@
In the last few years there has been some research in how to perform Bayesian optimization over combinatorial domains. As highlighted by [27], this task comes with two major challenges: constructing surrogate models that effectively capture the complexity of the combinatorial space, and efficiently searching within this space to identify the next candidate for evaluation using the acquisition function.
Most previous research has been devoted to the first task. Within this, an important line of investigation consists of modifying GP kernel functions to accommodate categorical covariate spaces. For instance [11] treat categorical variables as continuous ones via one-hot encoding; then, within the kernel function, continuous covariates are suitably transformed to one-hot encodings. [25] construct a combinatorial graph over the search space; subsequently, graph Fourier transforms are used to derive a diffusion kernel on the graph. Another interesting line of research performs BO over a continuous latent representation of the combinatorial spaces learned through a Variational Autoencoder [12]. Finally, a third line of research gets rid of GPs, and proposes using inherently discrete surrogate models. For instance, [15] use random forests to model the objective function given the covariates. [2] model the response as a linear function of the covariates including all possible pairwise interactions. A heavy-tailed horseshoe prior is used to induce sparsity on the coefficients.
@@
In this paper, we address the second key challenge in BO: efficiently searching within a combinatorial domain for the next candidate to evaluate after selecting an appropriate surrogate model. We introduce a novel simulation-based approach for optimizing acquisition functions in combinatorial domains, which we refer to as Simulation-Based Bayesian Optimization (SBBO). The key advantage of SBBO is its ability to optimize the acquisition function using only samples from the posterior predictive distribution, without requiring a closed-form expression. This makes BO applicable even when posterior inference is carried out through methods like Markov Chain Monte Carlo (MCMC). By removing the need for closed-form solutions, SBBO allows for the use of a broader range of surrogate probabilistic models that may be better suited for BO tasks in combinatorial spaces. This flexibility makes SBBO particularly effective in these contexts and significantly expands the applicability of BO.
@@ 2 Bayesian Optimization
Bayesian optimization (BO) is a class of methods centered on solving the optimization problem
@@
In the BO setting, the objective function $f(\boldsymbol{x})$ is assumed to be a black box, so no assumptions of, e.g., convexity or linearity are made. Additionally, it is assumed that $f$ is expensive to evaluate in that the number of evaluations of $f$ that can be feasibly performed is small, perhaps due to computing time or financial cost. It is further assumed that querying the function at $x$ provides a noisy observation $y$ of $f(x)$.
@@
BO uses a surrogate probabilistic model of the objective function, usually obtained through Bayesian methods, to inform the next evaluation decision. It works iteratively by: 1) fitting the surrogate model, and 2) using an acquisition function to obtain a new point $\boldsymbol{x}$ at which we evaluate or observe $f(\boldsymbol{x})$. The surrogate model is typically fitted using all available data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. After selecting the new evaluation location $\boldsymbol{x}_{n+1}$, the corresponding $y_{n+1}$ is obtained, and the surrogate model is updated to include the additional data point $(\boldsymbol{x}_{n+1},y_{n+1})$. Let us briefly review these two key components of Bayesian Optimization.
For the surrogate model, the following observational equation is assumed
[…]
@@
where $f(\boldsymbol{x})$ could be parametric or non-parametric. Given a new location $\boldsymbol{x}$ and appropriate choice of prior, the posterior predictive distribution on $f(\boldsymbol{x})$ is taken as the surrogate model. If a parametric model is chosen for $f$, then $f(\boldsymbol{x})$ will depend on parameters $\boldsymbol{\beta}$ and is denoted as $f_{\boldsymbol{\beta}}(\boldsymbol{x})$. A prior $\pi(\boldsymbol{\beta})$ for the parameters $\boldsymbol{\beta}$ needs to be specified. Coupled with a likelihood $\pi(y|\boldsymbol{x},\boldsymbol{\beta},\sigma^{2})$, after observing data $\mathcal{D}_{1:n}$, we obtain the posterior $\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$. Then, the surrogate model is
@@
$\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})\equiv\pi(f_{\boldsymbol{\beta}}(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})=\mathbb{E}_{\boldsymbol{\beta}|\mathcal{D}_{1:n}}\left[f_{\boldsymbol{\beta}}(\boldsymbol{x})\right]$ (2.3)
Note that, in most cases, the expectation in (2.3) does not have an analytical form. Typically, the only option is to generate samples from the surrogate model by first sampling $\beta\sim\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$, usually via MCMC, and then computing $f_{\boldsymbol{\beta}}(\boldsymbol{x})$.
@@
If we choose instead a non-parametric model for $f$, then a prior $\pi(f)$ over $f$ needs to be specified. Coupled with the likelihood $\pi(y|f,\sigma^{2})$, we obtain the posterior distribution $\pi(f|\mathcal{D}_{1:n})$ over $f$. Then for a given new location $\boldsymbol{x}$, the evaluation of the posterior at $\boldsymbol{x}$, $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, is taken as the surrogate model.
@@ 3.1 Main Idea
We aim to solve the optimization problem 2.1 with a limited number of function evaluations. Suppose we have access to data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$, possibly obtained from previous experiments, where $\boldsymbol{x}_{i}$ is a vector of covariates for observation $i$ and $y_{i}$ is derived from the observation model (2.2). Given a budget of at most $M$ additional function evaluations, our task is to determine the next evaluation points $\boldsymbol{x}_{n+1},\boldsymbol{x}_{n+2},\ldots,\boldsymbol{x}_{n+M}$. As introduced in Section 2, these points are usually selected in a greedy manner. Let $u(f(\boldsymbol{x}),\boldsymbol{x})$ be a non-negative and bounded utility function that quantifies the benefits or gains from evaluating the objective function at $\boldsymbol{x}$ and observing the corresponding $f(\boldsymbol{x})$. Following Bayesian decision-theoretic principles, the next evaluation should be performed at the location $\boldsymbol{x^{*}}$ that maximizes the expected utility (2.4).
To perform Bayesian optimization, we need to select a suitable Bayesian surrogate model and optimize (2.4). As mentioned earlier, except in a few special cases, the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ is generally only accessible through sampling, typically using MCMC methods. The specifics of the sampling process depend on the surrogate model employed (see Section 3.3 for examples). Assuming we can generate samples from $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, we can borrow ideas from Bayesian decision analysis and reframe the optimization of (2.4) as a simulation problem. Following [4], we define an augmented distribution $g$ that treats $\boldsymbol{x}$ as a random variable:
@@
the mode of this distribution also coincides with the optimal point at which we should make our next function evaluation. In particular, if $H$ is sufficiently large, the marginal distribution $g_{H}(\boldsymbol{x})$ tightly concentrates on the next $\boldsymbol{x}_{n+1}$, thus facilitating mode identification.
@@
[…]
@@
If the utility is non-negative and bounded, and the state space of the inhomogeneous Markov chain generated by the sampler is finite, [24] demonstrates that the chain is strongly ergodic when the update schedule for $H$ is logarithmic. Additionally, its stationary distribution is uniform over the set of $\boldsymbol{x}$ that maximize (2.4). In practice, when the state space is not finite, it will be necessary to discretize $v$ into a finite grid.
We have now transformed the optimization of (2.4) into a simulation problem. The objective is to solve this optimization without explicitly evaluating the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, relying only on samples from it. Therefore, we need to simulate from the augmented distribution (3.1) without requiring an analytical expression for the posterior. This is where a key observation becomes crucial. Within the MCMC in Algorithm (1), we propose using $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ as the proposal distribution for $\tilde{f}(\boldsymbol{x})_{h}$ for $h=1,\dots,H$, once a candidate $\tilde{\boldsymbol{x}}_{n+1}$ has been proposed. This approach ensures that the acceptance probability becomes independent of the posterior density, allowing us to simulate from (3.1) without evaluating the posterior directly. To see this more clearly: if the current state of the chain is $(\boldsymbol{x},v)$, where $v=\frac{1}{H}\sum_{h=1}^{H}\log u(f(\mathbf{x})_{h},\mathbf{x})$, the proposed new state is $(\tilde{\boldsymbol{x}},\tilde{v})$, where $\tilde{v}=\frac{1}{H}\sum_{h=1}^{H}\log u(\tilde{f}(\mathbf{x})_{h},\tilde{\boldsymbol{x}})$. With $q(\cdot|\boldsymbol{x})$ as the proposal distribution for $\boldsymbol{x}$, the acceptance ratio for each iteration of Algorithm 1 is:
@@ 3.2 Gibbs Sampler
A possible variation of Algorithm (1) entails using a Gibbs sampler to simulate from $g_{H}$ rather than a Metropolis-Hastings scheme for every $H$. Recall that $\boldsymbol{x}$ is a vector of $p$ covariates. The conditional distribution of the $s$-th element of $\boldsymbol{x}$ can be written as
@@
For the particular case in which $\boldsymbol{x}$ is a vector of $p$ categorical variables, this marginal is simply the softmax distribution over $\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s})\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]$. However, sampling from this distribution requires evaluating posterior predictive probabilities, which defeats the purpose of SBBO. As an alternative, we could create a Gibbs sampling scheme to sample from $g_{H}(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{H}|\mathcal{D}_{1:n})$ where $(f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H})$ has been integrated out. Sampling from $g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s},,\mathcal{D}_{1:n})$ can be done using Metropolis steps. Assuming that the current state of the chain is $\boldsymbol{x}$, $f(\boldsymbol{x})_{1},\dots,f(\boldsymbol{x})_{H}$ and $v=\frac{1}{H}\sum_{h=1}^{H}\log u(f(\boldsymbol{x})_{h},\boldsymbol{x})$, we proceed as follows:

### v3

@@ 1 Introduction
Bayesian Optimization (BO, [23]) stands as an efficient approach for identifying the extrema of a black-box objective function. Its effectiveness originates from its capacity to integrate prior knowledge about the function with the evaluations previously conducted, enabling informed decisions on where to carry out additional evaluations. BO involves the construction of a Bayesian surrogate model for $f(\boldsymbol{x})$. Subsequently, evaluation points are selected through the maximization of an acquisition function [11].
@@
In the last few years, there has been growing interest in performing Bayesian optimization over combinatorial domains. As highlighted by [32], this task presents two major challenges: (i) constructing surrogate models that capture the complexity of the combinatorial space and (ii) efficiently searching that space to select the next point for evaluation. Most previous research has focused on the first challenge, often by retaining GP surrogates while designing specialized kernel functions adapted to discrete structures. In these approaches, the GP’s covariance function is modified to reflect similarity between discrete configurations rather than Euclidean distance. For example, [13] simply treat categorical (encoded via one-hot) and integer-valued variables as continuous and apply rounding before including them in the kernel. Other works rely on custom kernels based on Hamming distance or similar metrics (e.g., treating two configurations as more similar if they share more categorical components), thus aligning better with the discrete nature of the domain. Some methods employ string kernels (e.g., substring matching kernels) to model sequence data; for instance, [24] propose “BOSS”, a GP-based BO method that uses a subsequence string kernel to avoid naive one-hot encodings. Similarly, [28] develop a GP with a diffusion kernel on a graph representing the combinatorial space. In their COMBO approach, each discrete value is treated as a vertex in a graph, and a diffusion-based covariance is defined over these vertices, enabling BO in high-dimensional binary and categorical inputs while preserving GP-based uncertainty. For an in-depth review of BO over discrete sequences, we refer the reader to [14].
Beyond using fixed kernels, some methods combine GPs with learned embeddings or hybrid kernels to address more complex discrete domains. Instead of defining a kernel directly in the original combinatorial space, one can map discrete inputs into a learned continuous representation and then apply standard GP kernels. [7] follow this approach by using an autoencoder to embed sequences in a latent space, combining a latent-space kernel with a structured kernel in the original space. This hybrid kernel strategy enables the GP to capture both high-level features (through the latent mapping) and the finer discrete structure (through a dedicated combinatorial kernel). More recently, [8] propose a Hamming embedding via dictionaries, which preserves Hamming distances in a continuous space so that RBF or Matérn kernels can still reflect categorical similarity. Overall, a rich body of work has focused on retaining GPs while adapting their kernels to discrete search spaces.
An alternative research direction replaces GPs with surrogates that more naturally accommodate high-dimensional or categorical inputs. One classical example is the use of random forests, as popularized by SMAC [17], which discards the full GP posterior but scales more easily to many or purely categorical variables. Another category of GP alternatives uses parametric models, such as Bayesian linear regressions. For example, the proposal by [2], named BOCS (Bayesian Optimization of Combinatorial Structures), replaces the GP with a sparse Bayesian linear regression over polynomial expansions of binary inputs, placing a horseshoe prior on coefficients to identify a small subset of important interaction terms. This approach can be especially effective when the black-box function has mostly low-order interactions. Researchers have also investigated Bayesian neural network surrogates for high-dimensional BO, as in BOHAMIANN [31], where a Bayesian neural network serves as the surrogate.
@@
In this paper, rather than designing a new surrogate model, we address the second challenge in BO for combinatorial spaces: efficiently searching a discrete domain for the next candidate once a suitable surrogate has been chosen. We introduce a novel simulation-based approach for optimizing acquisition functions in combinatorial domains, which we refer to as Simulation-Based Bayesian Optimization (SBBO). The key advantage of SBBO is its ability to optimize the acquisition function using only samples from the posterior predictive distribution, without requiring a closed-form expression. This makes BO applicable even when posterior inference is carried out through methods such as Markov Chain Monte Carlo (MCMC). By dispensing with closed-form assumptions, SBBO can leverage broader classes of surrogate probabilistic models that may be better aligned with the demands of combinatorial BO. This flexibility is particularly beneficial in high-dimensional discrete problems, where GPs are often criticized for poor scalability and the difficulty in defining meaningful kernel similarities as dimension increases.
@@ 2 Bayesian Optimization
Bayesian optimization is a class of methods centered on solving the optimization problem
@@
In the BO setting, the objective function $f(\boldsymbol{x})$ is assumed to be a black box, so no assumptions of convexity or linearity, for example, are made. Additionally, it is assumed that $f$ is expensive to evaluate in that the number of evaluations of $f$ that can be feasibly performed is small, perhaps due to computing time or financial cost. It is further assumed that querying the function at $x$ provides a noisy observation $y$ of $f(x)$.
@@
BO uses a surrogate probabilistic model of the objective function, usually obtained through Bayesian methods, to inform the next evaluation decision. It works iteratively by: 1) fitting the surrogate model, and 2) using an acquisition function to obtain a new point $\boldsymbol{x}$ at which we evaluate or observe $f(\boldsymbol{x})$. The surrogate model is typically fitted using all available data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. After selecting the new evaluation location $\boldsymbol{x}_{n+1}$, the corresponding $y_{n+1}$ is obtained and the surrogate model is updated to include the additional data point $(\boldsymbol{x}_{n+1},y_{n+1})$. We now briefly review these two key components of Bayesian Optimization.
For the surrogate model, the following observational equation is assumed:
$y\equiv y(\boldsymbol{x})=f(\boldsymbol{x})+\epsilon,~~~~~\epsilon\overset{\text{iid}}{\sim}\mathcal{N}(0,\sigma^{2})$ (2.2)
@@
where $f(\boldsymbol{x})$ could be parametric or non-parametric. Given a new location $\boldsymbol{x}$ and appropriate choice of prior, the posterior predictive distribution on $f(\boldsymbol{x})$ is taken as the surrogate model. If a parametric model is chosen for $f$, then $f(\boldsymbol{x})$ will depend on parameters $\boldsymbol{\beta}$ and is denoted as $f(\boldsymbol{x})\equiv f_{\boldsymbol{\beta}}(\boldsymbol{x})$. A prior $\pi(\boldsymbol{\beta})$ must be specified for $\boldsymbol{\beta}$, which, when combined with the likelihood $\pi(y|\boldsymbol{x},\boldsymbol{\beta},\sigma^{2})$, yields the posterior $\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$ after observing data $\mathcal{D}_{1:n}$. The surrogate model is then the push-forward of this posterior through the mapping $\boldsymbol{\beta}\mapsto f_{\boldsymbol{\beta}}(\boldsymbol{x})$:
@@
$\pi(f(\boldsymbol{x})=z|\boldsymbol{x},\mathcal{D}_{1:n})\equiv\int\delta\left(z-f_{\boldsymbol{\beta}}(\boldsymbol{x})\right)\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})\mathop{}\!\mathrm{d}\boldsymbol{\beta}$ (2.3)
In most cases, the integral (2.3) lacks a closed-form solution. In these scenarios, surrogate model samples are typically generated by first drawing $\boldsymbol{\beta}\sim\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$, often via MCMC, and then evaluating $f_{\boldsymbol{\beta}}(\boldsymbol{x})$.
@@
If we instead choose a non-parametric model for $f$, then a prior $\pi(f)$ over $f$ needs to be specified. Coupled with the likelihood $\pi(y|f,\sigma^{2})$, we obtain the posterior distribution $\pi(f|\mathcal{D}_{1:n})$ over $f$. Then for a given new location $\boldsymbol{x}$, the evaluation of the posterior at $\boldsymbol{x}$, $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, is taken as the surrogate model.
@@ 3.1 Main Idea
We aim to solve the optimization problem (2.1) with a limited number of function evaluations. Suppose we have access to data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$, possibly obtained from previous experiments, where $\boldsymbol{x}_{i}$ is a vector of covariates for observation $i$ and $y_{i}$ is derived from the observation model (2.2). Given a budget of at most $M$ additional function evaluations, our task is to determine the next evaluation points $\boldsymbol{x}_{n+1},\boldsymbol{x}_{n+2},\ldots,\boldsymbol{x}_{n+M}$. As introduced in Section 2, these points are usually selected in a greedy manner. Following Bayesian decision-theoretic principles, the next evaluation should be performed at the location $\boldsymbol{x^{*}}$ that maximizes the expected utility (2.4).
To perform Bayesian Optimization, we need to select a suitable Bayesian surrogate model and optimize (2.4). As mentioned earlier, except in a few special cases, the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ is generally only accessible through sampling, typically using MCMC methods. When only sampling access to the posterior predictive distribution is available, a straightforward approach to optimizing (2.4) would be to construct a Monte Carlo (MC) approximation of the expected utility and optimize this approximation directly. However, this approach encounters significant difficulties. Because the expectation defining the acquisition function depends explicitly on the covariates $\boldsymbol{x}$, a separate MC approximation must be computed for every candidate solution explored during optimization, substantially increasing computational complexity. Moreover, as noted by [27], the expected utility surfaces tend to be particularly flat around their maxima, especially in high-dimensional scenarios. Consequently, MC approximation errors can obscure the true optimal regions, making them challenging to identify without an impractically large number of samples.
To address these limitations, we propose an alternative methodology drawn from the Bayesian decision analysis literature. Assuming we can generate [^1] The specifics of the sampling process depend on the surrogate model employed (see Section 3.3 for examples). samples from $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, and that the utility function $u(f(\boldsymbol{x}),\boldsymbol{x})$ is positive and bounded222 Note that this positivity constraint is mild, as any bounded utility function can be shifted and scaled to become strictly positive without affecting the optimal decisions.; we can borrow ideas from Bayesian decision analysis and reframe the optimization of (2.4) as a simulation problem. Following [4], we define an augmented distribution $g$ that treats $\boldsymbol{x}$ as a random variable:
@@
the mode of this distribution also coincides with the optimal point at which we should make our next function evaluation. In particular, if $H$ is sufficiently large, the marginal distribution $g_{H}(\boldsymbol{x})$ tightly concentrates on the next $\boldsymbol{x}_{n+1}$, thus facilitating mode identification. Using this sharpened target is key to overcoming the challenges of direct optimization. When the acquisition function is relatively flat, naive Monte Carlo approximations are often ineffective, as stochastic errors can obscure the true peaks and cause the search to fail. In contrast, by sampling from a sharper version of the expected utility, our method concentrates probability density around the function’s mode, providing a much more robust search.
@@
    $(\boldsymbol{x}_{n+1}^{(1)},v^{(1)})=\begin{cases}(\tilde{\boldsymbol{x}}_{n+1},\tilde{v})&w.p.\ \alpha\\ (\boldsymbol{x}^{(0)}_{n+1},v^{(0)})&w.p.\ 1-\alpha\end{cases}$
@@
If the utility is positive and bounded, and the state space of the inhomogeneous Markov chain generated by the sampler is finite, [27] demonstrate that the chain is strongly ergodic when the update schedule for $H$ is logarithmic. Additionally, its stationary distribution is uniform over the set of $\boldsymbol{x}$ that maximize (2.4). In practical situations where the state space is continuous or infinite, discretization of $v$ onto a finite grid becomes necessary.
In principle, as $H$ increases, the marginal distribution over $\boldsymbol{x}$ concentrates sharply around the global mode. Therefore, taking the last visited state of Algorithm 1 could theoretically yield the optimal evaluation point. However, in practice, to enhance numerical stability, we heuristically approximate the global mode by independently estimating the modal value for each covariate dimension from the simulated samples.
[…]
@@ 3.2 Gibbs Sampler
A possible variation of Algorithm 1 entails using a Gibbs sampler to simulate from $g_{H}$ rather than a Metropolis-Hastings scheme for every $H$. Recall that $\boldsymbol{x}$ is a vector of $p$ covariates. The conditional distribution of the $s$-th element of $\boldsymbol{x}$ can be written as
@@
For the particular case in which $\boldsymbol{x}$ is a vector of $p$ categorical variables, this marginal is simply the softmax distribution over $\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s})\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]$. However, sampling from this distribution requires evaluating posterior predictive probabilities, which defeats the purpose of SBBO. The Gibbs sampler is thus beneficial only for scenarios in which the predictive distribution has a closed-form expression but the acquisition function does not. In such cases, the Gibbs sampler leverages the known predictive distribution to yield closed-form conditional marginals in the covariate space, significantly accelerating computation compared to a full Metropolis-Hastings scheme.
When closed-form predictive distributions are not available, we alternatively could create a Gibbs sampling scheme to sample from $g_{H}(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{H}|\mathcal{D}_{1:n})$ where $(f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H})$ has been integrated out. Sampling from $g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s},,\mathcal{D}_{1:n})$ can be done using Metropolis steps. Assuming that the current state of the chain is $\boldsymbol{x}$, $f(\boldsymbol{x})_{1},\dots,f(\boldsymbol{x})_{H}$ and $v=\frac{1}{H}\sum_{h=1}^{H}\log u(f(\boldsymbol{x})_{h},\boldsymbol{x})$, we proceed as follows:
@@
Another possible variation aside from the Gibbs sampler is to use Sequential Monte Carlo (SMC) methods to approximate both the posterior distribution and the expected utility. For instance, [29] proposed employing SMC to sequentially approximate the posterior over the parameters of a surrogate parametric model in Bayesian optimization, subsequently computing expected utilities as weighted sums over the resulting particles. In our context, however, an even more suitable approach could be the interacting particle system proposed by [1], which samples from progressively annealed versions of the augmented distribution (3.1), gradually concentrating particles around its mode. Adapting this sampling strategy to Bayesian optimization represents an intriguing direction for extending the SBBO framework.